# Automatically Detecting Checked-In Secrets in Android Apps: How Far Are We?

### v1

@@ Abstract.
Mobile apps are predominantly integrated with cloud services to benefit from enhanced functionalities. Adopting authentication using secrets such as API keys is crucial to ensure secure mobile-cloud interactions. However, developers often overlook the proper storage of such secrets, opting to put them directly into their projects. These secrets are checked into the projects and can be easily extracted and exploited by malicious adversaries. While many researchers investigated the issue of checked-in secret in open-source projects, there is a notable research gap concerning checked-in secrets in Android apps deployed on platforms such as Google Play Store. Unlike open-source projects, the lack of direct access to the source code and the presence of obfuscation complicates the checked-in secret detection for Android apps. This motivates us to conduct an empirical analysis to measure and compare the performance of different checked-in secret detection tools on Android apps. We first conducted a literature review to find all the checked-in secret detection tools that can be applied to Android apps. Then, we evaluate three representative tools on 5,135 Android apps, comparing their performance and analyzing their limitations. Our experiment reveals 2,142 checked-in secrets affecting 2,115 Android apps. We also disclose that the current checked-in secret detection techniques suffer from key limitations.
All of the evaluated tools can miss a significant number of checked-in secrets in Android apps.
@@
Nevertheless, we observed that the tools are complimentary, suggesting the possibility of developing a more effective checked-in secret detection tool by combining their insights. Additionally, we propose that analyzing string groups within methods containing checked-in secrets may provide a more effective strategy to overcome obfuscation challenges.
Android apps, Checked-in Secret, Empirical Study
ccs: Security and Privacy Detections of Leaked Checked-in Secrets
1. Introduction
Cloud computing is one of the most prevalent technologies in the modern software industry (Michalowski, [n. d.]).
Cloud services typically require clients to use secrets (e.g., API keys or API secrets) for authentication and secure access. However, in practice, developers often neglect the protection of the secrets and check them into their projects (Basak et al., 2023c).
For instance, the usage of hard-coded secrets is one of the 25 most dangerous software weaknesses (CWE, 2023). Such checked-in secrets can be exploited once extracted from the released software projects (798, 2024).
As an example, malicious users can use the extracted secrets to gain access to sensitive information. Recently, the database secrets of Uber company were mistakenly uploaded to GitHub, resulting in a leakage of information of 50,000 Uber drivers (ube, 2016).
@@ 1. Introduction
Most of the existing studies focused on analyzing the checked-in secret problem in code-sharing platforms such as Github or Gitlab. For example, Meli et al. (Meli et al., 2019) designed a framework based on regular expressions that can identify potential checked-in secrets within a GitHub repository. Feng et al. (Feng et al., 2022) implemented a machine learning tool, PassFinder, to predict whether a string is a checked-in secret based on its contextual surroundings and intrinsic value. There also exist a number of community tools that can detect checked-in secrets in code-sharing platforms (tru, 2024; ggs, 2024; git, 2022; rep, 2024; whi, 2024; gsh, 2024). Other than developing a checked-in secret detection tool, some studies (Basak et al., 2023c; Rahman et al., 2022; Krause et al., 2022) conducted empirical analyses to investigate the checked-in secret issue on the code-sharing platform.
@@
Meanwhile, there is a noticeable research gap of the checked-in secret problem on Android apps that are deployed on the Google Play Store. Android apps are also known to be susceptible to the problem of checked-in secrets (Basak et al., 2023c). Checked-in secrets can be extracted from the deployed Android apps through unpacking and reverse engineering (Renato, [n. d.]). Nevertheless, few studies have investigated this problem. Zuo et al. (Zuo et al., 2019) proposed the only tool, LeakScope, for checked-in secret detection in Android apps. Their study revealed that over 15,000 Android apps have the checked-in secret problem. However, their study is limited to three cloud providers, Google, Amazon, and Microsoft and requires substantial time and computational resources to expand its detection capabilities and analyze Android apps.
@@
In this paper, we aim to evaluate the state-of-the-art checked-in secret detection tools when applied in Android apps and provide insights for the future development of Android checked-in secret detectors. We evaluate the performance of Android-specific checked-in secret detector, LeakScope, as well as adapting source-code-level checked-in secret detectors designed for code-sharing platforms. The source-code-level checked-in secret detectors can be adapted to Android apps since the source code of the apps can be recovered through decompiling or reverse engineering. Meanwhile, it is still challenging to directly leverage the source-code level checked-in secret detectors in Android apps due to the wide adoption of obfuscation. Obfuscation obscures the lexical information of the code that is often needed by the source-code-level checked-in secret detectors.
@@
We aim to answer the following research questions in our paper:
@@
  • RQ1: (Survey of existing tools) What are the existing checked-in secret detection tools? Which tools can be adapted to Android apps?
@@
  • RQ3: (Implications) What are the limitations of the evaluated tools? What are the major implications of our research?
To address the proposed research questions, we conducted a systematic literature review to identify all academic and community checked-in secret detection tools. We discovered seven academic checked-in secret detection tools and six community tools. We further categorized the identified tools into three categories based on their methodology: Intrinsic value analysis, static analysis and machine learning-based analysis. We selected the framework proposed by Meli et al. (intrinsic value)  (Meli et al., 2019), LeakScope (static)  (Zuo et al., 2019), and PassFinder (machine learning)  (Feng et al., 2022) as the representative tool for each category.
@@
  • We provided a benchmark dataset consisting of the code snippets of the validated checked-in secrets in Android apps.
Data availability. Our code and dataset are available on our project website (dat, [n. d.]). For the dataset, we provided sanitized code snippets and data flow of the detected checked-in secrets. Due to the sensitiveness of the dataset, we did not provide the original Android apps we used in the experiment. The Android apps can be shared to selected researchers and developers upon request.
2. Background
2.1. Checked-in Secrets
@@
Cloud service providers deploy various methods for authenticating processes to minimize security risks (Lu, 2011). Using the secrets to authenticate cloud service is the most common approach. When incorporating cloud services into an application, cloud service providers require developers to generate secrets.
@@
These secrets allow the cloud service providers to identify which project is calling the cloud API, verify its access to certain resources, and manage billing functions.
@@
Leaking the sensitive cloud service secrets to the public can result in severe security consequences for both cloud service providers and applications. Access to the checked-in secrets allows anyone to utilize the allocated resources and functionalities. It can lead to (1) data leakage if the checked-in secret allows for access to application data, (2) data corruption if the checked-in secret allows for data manipulation, (3) monetary losses if it creates costs for invoking the concerned APIs, and (4) denial of service (DoS) if there is a limit for the API invocation.
{listing}
@@
Code snippet with one checked-in secret. It is extracted from an Android app in our experiments.
{minted}
[linenos, frame=single, obeytabs=true, numbersep=-10pt]java task.addOnSuccessListener(listener) context = r0.getApplicationContext()¿() r7.ggh(context, ”AIza——”) r0.setupLifecycleListener()
@@
As an example, listing 1 is a code snippet extracted from one Android app. The string at line 3 that starts with ”Alza” is a Google API key, which could be used to grant access to Google Map services. This secret can be used to perform a DoS attack since the API has a limit of requests per minute or cause financial losses since it costs \$0.005 per request.
2.2. APK Files
@@
Unlike the open-source Android app projects on GitHub, the Android apps deployed on the Google Play Store are distributed as APK (Android Package Kit) files. An APK file is essentially an archive file that contains all the information needed for an Android app to be successfully executed on a device. When a user wants to download an Android app from platforms such as the Google Play Store, the platforms will automatically download the APK files for the users. Several automatic tools exist, such as ApkTool (apk, 2024), dex2jar (dex, 2024), Soot (soo, 2024) and Mobsf (mob, 2024), that can perform reverse engineering on an APK file. However, these reverse engineering tools all have limitations. They may not accurately recover the source code from APK files. Sometimes, they might even miss some classes while analyzing the APK files. Furthermore, APK files commonly adopt obfuscation techniques to enhance the application’s security (Dong et al., 2018). Obfuscation obscures the code of the Android apps by altering the names and hierarchies of code artifacts. For example, in line 3 of Listing  4.1.4, the method ggh was originally a method called ”initialize()” provided by Google to send an API request. Those factors are the main challenges that need to be addressed for the tools focused on Android apps. The effectiveness of detection tools relying on code context is significantly reduced by obfuscation, which makes variables and class names meaningless.
3. RQ1: Survey of Existing Tools
@@
The checked-in secrets issue is a long-studied subject in the software development field. There exist many tools that aim to detect checked-in secret. In RQ1, we aim to find all these available academic and community checked-in secret detection tools and analyze whether they can be adopted on Android apps through a detailed literature review process.
3.1. RQ1-Methodology: Literature Review
@@
We conducted a literature review to gather a comprehensive list of checked-in secret detection tools/methodologies. Our search query is composed of two parts, as shown in Table 1: Part 1 includes terms related to secret types (e.g., API key, secret, password), while Part 2 focuses on detection methodologies.
Table 1. Search String Group
@@
Each search query is a combination of Part 1 and Part 2, resulting in a total of 12 queries (i.e., 4*3), such as ”API key detection.” We initially searched these keywords across Google Scholar, IEEE Xplore, and the ACM Digital Library.
@@
Each query returns a large number of search results, including irrelevant ones. For example, some queries, such as API key detection, returned over 10,000 results on ACM digital library. Manually analyzing and selecting relevant papers from such a large pool is impractical. Therefore, for each search query, only the first 500 results were analyzed. In total, 5,500 papers were selected for our initial filtering process.
@@
During the initial filtering process, any paper with a title not related to terms like API, security, secret detection, or privacy leakage was excluded. This step resulted in the selection of 78 papers. Subsequently, two authors conducted a more thorough review of these 78 papers to identify those most relevant to our subject. In this stage, both authors read the abstracts (and, if necessary, the full content) of each paper to determine its relevance. If there was any disagreement, a discussion was held to reach a consensus. 19 papers were selected after this step.
@@
We also conducted a snowballing process, a strategy that is commonly adopted in other literature review research (Liu et al., 2022). This ensures we do not miss any important papers on the checked-in secret. For each paper that we initially selected, we reviewed references from the selected papers as well as papers that cited them. This process continued until no new relevant papers were identified.
@@
Eventually, we found 29 academic papers that are related to checked-in secret detection.
@@
For community tools, we used the six community tools selected by Setu et al.  (Basak et al., 2023a) as they conducted a thorough selection process to perform an analysis of checked-in secret detection tools on the open-source repositories.
3.2. RQ1-Result
@@
Among the 29 papers, seven of them developed a checked-in secret detection tool or methodology of our interest, while others are either empirical studies on checked-in secret issue and management or only specific to platforms such as DockerHub. Overall, a total of 13 checked-in secrets detection tools when combined with community-developed tools.
@@
We categorized the checked-in secret detection tools into three groups: Intrinsic value analysis, static analysis and Machine Learning based analysis according to their detection approach. Table 2 presents the list of all the identified checked-in secret detection tools and their corresponding categorization. Overall, there are seven Intrinsic value analysis tools, one static analysis tool and five Machine Learning based analysis tools.
@@
Intrinsic value analysis. This type of checked-in secret detection tools focused on the intrinsic value of the strings under analysis. The analysis is generally done using two processes. First, regular expression matching. In the cases of certain cloud providers, the generated checked-in secrets often adhere to the same format. For instance, Google’s checked-in secrets always take the form of “ALza******”. Utilizing regular expression matching enables the extraction of such checked-in secrets from the code. Second, entropy calculation. As checked-in secrets are all randomly generated strings, the randomness of the characters in a string will be a good indicator of checked-in secret. These techniques leveraged Shannon entropy to give a good estimation of randomness. A randomly generated string will generally have higher entropy than a normal string.
Static analysis.
@@
To the best of our knowledge, there has only been one tool that extracts checked-in secrets through static analysis, which is LeakScope(Zuo et al., 2019). By providing a location where a checked-in secret is used, static analysis can calculate the value of the checked-in secret through data and control flow analysis.
@@
Machine Learning based analysis. This type of checked-in secret detection tool takes into consideration the fact that the contextual surroundings of checked-in secrets follow some specific patterns. Factors such as file type, class name, variable name and code snippet could be used as features to train a machine learning model to predict whether a string is a checked-in secret.
Table 2. Categorization and Tools
Category Tools Seletced
Intrinsic Value Analysis Meli et al (Meli et al., 2019)
Gitleak (git, 2022)
Repo Supervisor (rep, 2024)
Trufflehog (tru, 2024)
Whispers (whi, 2024)
GGshield (ggs, 2024)
Github Secret Scanner (gsh, 2024)
Static Analysis LeakScope (Zuo et al., 2019)
Machine Learning Analysis PassFinder (Feng et al., 2022)
Secrethunter (Wen et al., 2022)
Seagull  (Han et al., 2024)
Saha et al (Saha et al., 2020)
Lounici et al (Lounici et al., 2021)
3.3. Tool Selection for Further Evaluation.
The checked-in secret detection tools within the same category adopted similar techniques. For example, checked-in secret detection tools of type intrinsic value analysis either adopted regular expression matching or entropy calculation. Some tools adopted both of the two techniques. Therefore, we decided to select one representative tool for each category for further evaluation in RQ2 and RQ3.
For intrinsic value analysis, we chose the Three-Layer Filter technique proposed by Meli et al. (Meli et al., 2019) since (1) intrinsic value analysis highly relies on manually crafted regular expressions; the Three-Layer Filters provided a comprehensive list of precise regular expressions, and (2) the Three-Layer Filter is proven to be one of the best tools in this category (Basak et al., 2023a; Meli et al., 2019).

### v2

@@ Abstract
Mobile apps are predominantly integrated with cloud services to benefit from enhanced functionalities. Adopting authentication using secrets such as API keys is crucial to ensure secure mobile-cloud interactions. However, developers often overlook the proper storage of such secrets, opting to put them directly into their projects. These secrets are checked into the projects and can be easily extracted and exploited by malicious adversaries. While many researchers investigated the issue of checked-in secret in open-source projects, there is a notable research gap concerning checked-in secrets in Android apps deployed on platforms such as Google Play Store. Unlike open-source projects, the lack of direct access to the source code and the presence of obfuscation complicates the checked-in secret detection for Android apps. This motivates us to conduct an empirical analysis to measure and compare the performance of different checked-in secret detection tools on Android apps.
We conducted a systematic literature review to identify existing checked-in secret detection tools and then analyzed how these tools could be adapted and applied to the context of Android apps. Then, we evaluated three representative tools: Three-Layer Filter, LeakScope, and PassFinder on 5,135 Android apps, comparing their performance and analyzing their limitations. Our experiment reveals 2,142 checked-in secrets affecting 2,115 Android apps and also discloses that the current checked-in secret detection techniques suffer from key limitations. While all the tools achieved high precision, each of them can miss a significant number of checked-in secrets in Android apps.
@@
Nevertheless, we observed that the tools are complimentary, suggesting the possibility of developing a more effective checked-in secret detection tool by combining their insights. Additionally, we outline potential directions for future research, supported by preliminary experimental results.
1 Introduction
Cloud computing is one of the most prevalent technologies in the modern software industry cstat. Cloud services typically require clients to use secrets (e.g., API keys or API secrets) for authentication and secure access. In this study, we focus on such cloud API keys that are hardcoded in Android applications as prior work has shown they are the most prevalent type of leaked secret in practice twitterregex. Because, in practice, developers often neglect the protection of the secrets and check them into their projects basak2023challenges. For instance, the usage of hard-coded secrets is one of the 25 most dangerous software weaknesses CWE. Such checked-in secrets can be exploited once extracted from the released software projects 798. As an example, malicious users can use the extracted secrets to gain access to sensitive information. In 2017, the API keys of Uber company were mistakenly uploaded to GitHub, resulting in a leakage of information of 50,000 Uber drivers uber.
@@ 1 Introduction
Most of the existing studies focused on analyzing the checked-in secret problem in code-sharing platforms such as Github or Gitlab. For example, Meli et al. meli designed a framework based on regular expressions that can identify potential checked-in secrets within a GitHub repository. Feng et al. passfinder implemented a machine learning tool, PassFinder, to predict whether a string is a checked-in secret based on its contextual surroundings and intrinsic value. There also exist a number of community tools that can detect checked-in secrets in code-sharing platforms trufflehog; ggshield; gitleaks; repo-supervisor; whisper; gsh. Other than developing a checked-in secret detection tool, some studies basak2023challenges; rahman; krause2022committed conducted empirical analyses to investigate the checked-in secret issue on the code-sharing platform.
@@
Meanwhile, there is a noticeable research gap of the checked-in secret problem on Android apps deployed on the Google Play Store. Android apps are also known to be susceptible to the problem of checked-in secrets basak2023challenges. Checked-in secrets can be extracted from the deployed Android apps through unpacking and reverse engineering extractTutorial. Nevertheless, few studies have investigated this problem. Zuo et al. LeakScope proposed the only tool, LeakScope, for checked-in secret detection in Android apps. Their study revealed that over 15,000 Android apps have the checked-in secret problem. However, their study is limited to three cloud providers, Google, Amazon, and Microsoft and requires substantial time and computational resources to expand its detection capabilities and analyze Android apps. While this research aims to broaden the scope of checked-in secret detection, it also highlights a critical challenge: many Android applications available on app stores are only distributed as compiled APKs and are frequently obfuscated obfuscationprac. This combination restricts access to high-level code constructs and undermines techniques that depend on contextual or semantic information, such as machine-learning-based approaches. Additionally, since APKs are also the only artifacts accessible to potential attackers, evaluating secret detection in this format realistically simulates how secrets may be retrieved from closed-source apps. Furthermore, because the Google Play Store is the official marketplace and central distribution channel in the Android ecosystem, studying apps in this setting is both practically significant and security-relevant. These characteristics make Android apps particularly difficult to analyze and highlight why dedicated studies of secret detection in this setting are necessary.
@@
In this paper, we aim to fill the existing research gap by evaluating the performance of state-of-the-art checked-in secret detection tools, each representing a distinct detection approach, when applied to Android apps. It also provides insights for the future development of Android checked-in secret detectors. We evaluate the performance of Android-specific checked-in secret detector, LeakScope, as well as adapting source-code-level checked-in secret detectors designed for code-sharing platforms. The source-code-level checked-in secret detectors can potentially be adapted to Android apps since the source code of the apps can be recovered through decompiling or reverse engineering. However, it is still challenging to directly leverage the source-code level checked-in secret detectors in Android apps due to the wide adoption of obfuscation. Obfuscation obscures the lexical information of the code that is often needed by the source-code-level checked-in secret detectors. We also want to study the impact of obfuscation on these source-code-level checked-in secrets detectors.
@@
In summary, we want to answer the following research questions in this work:
@@
  • RQ1: (Survey of existing tools) What are the existing checked-in secret detection approaches? Which tools can be adapted to Android apps?
@@
  • RQ3: (Implications) What are the limitations of the evaluated tools? How can we improve checked-in secret detection for Android apps?
To achieve the proposed research objectives, we conducted a systematic literature review to identify the available academic and community checked-in secret detection tools. we discovered seven academic checked-in secret detection tools and six community tools. we further categorized the identified tools into three categories based on their methodology: Intrinsic value analysis, static analysis and machine learning-based analysis. we selected the framework proposed by Meli et al. (intrinsic value)  meli, LeakScope (static analysis)  LeakScope, and PassFinder (machine learning)  passfinder as the representative tool for each category based on their performance, availability, and adaptability.
@@
  • We provided a benchmark dataset consisting of the sanitized code snippets of the validated checked-in secrets in Android apps and the complete dataset can be access upon request.
2 Background
2.1 Checked-in Secrets
@@
Cloud service providers deploy various methods for authenticating processes to minimize security risks lu2011accessing. Using the secrets to authenticate cloud service is the most common approach. When incorporating cloud services into an application, cloud service providers require developers to generate secrets.
@@
These secrets allow the cloud service providers to identify which project is calling the cloud API, verify whether it has access to certain resources, and manage billing functions.
@@
Leaking the sensitive cloud service secrets to the public can result in severe security consequences for both cloud service providers and applications. Access to the checked-in secrets allows anyone to utilize the allocated resources and functionalities. It can lead to (1) data leakage if the checked-in secret is used to authenticate access to application data, (2) data corruption if the checked-in secret allows for data manipulation, (3) monetary losses if it creates costs for invoking the concerned APIs, and (4) denial of service (DoS) if there is a limit for the API invocation.
@@
Listing 1: Code snippet with one checked-in secret. It is extracted from an Android app in our experiments.
1task.addOnSuccessListener(listener)
2context = r0.getApplicationContext()>()
3r7.ggh(context, "AIza------")
4r0.setupLifecycleListener()
@@
As an example, Listing 1 is a code snippet extracted from one Android app in our experiment dataset. The string at line 3 that starts with “Alza”is a Google API key, which could be used to grant access to Google Map services. This secret can be used to perform a DoS attack since the API has a limit of requests per minute or cause financial losses since it costs \$0.005 per request.
2.2 APK Files
@@
Unlike the open-source Android app projects on GitHub, the Android apps deployed on the Google Play Store are distributed as APK (Android Package Kit) files. An APK file is essentially an archive file that contains all the information needed for an Android app to be successfully executed on a device. When a user wants to download an Android app from platforms such as the Google Play Store, the platforms will automatically download the APK files for the users. Several automatic tools exist, such as ApkTool apktool, dex2jar dex2jar, Soot soot, and Mobsf mobsf, that can perform reverse engineering on an APK file. However, these reverse engineering tools all have limitations. They may not accurately recover the source code from APK files. Sometimes, they might even miss some classes while analyzing the APK files. Furthermore, APK files commonly adopt obfuscation techniques to enhance the application’s security obfuscation. Obfuscation obscures the code of the Android apps by altering the names and hierarchies of code artifacts. For example, in line 3 of Listing 1, the method ggh was originally a method called “initialize()” provided by Google to send an API request. Those factors are the main challenges that need to be addressed for the tools focused on Android apps. The effectiveness of detection tools relying on code context is significantly reduced by obfuscation, which makes variables and class names meaningless.
3 RQ1: Survey of Existing Tools
@@
The checked-in secrets issue is a long-studied subject in the software development field. There exist many tools that aim to detect checked-in secret. In RQ1, we aim to find these available academic and community checked-in secret detection tools and analyze whether they can be adopted on Android apps through a detailed literature review process.
3.1 RQ1-Methodology: Literature Review
@@
We conducted a literature review to gather a comprehensive list of checked-in secret detection tools/methodologies. The search query is composed of two parts, as shown in Table 1: Part 1 includes terms related to secret types (e.g., API key, secret, password), while Part 2 focuses on detection methodologies.
Table 1: Search String Group
@@
Each search query combined Part 1 and Part 2, yielding 12 queries (4 × 3), such as “API key detection.” We conducted the initial search in Google Scholar, IEEE Xplore, and the ACM Digital Library in January 2024.
@@
Each query returns a large number of search results, including irrelevant ones. For example, some queries, such as API key detection, returned over 10,000 results on the ACM digital library. Manually analyzing and selecting relevant papers from such a large pool is impractical. Therefore, for each search platform, we chose to sort the results by relevance, as this is typically the default and recommended ordering used by digital libraries to display the most pertinent publications first. This approach increases the likelihood of capturing high-quality and topically appropriate papers early in the result set. And if a query returned more than 500 results, only the top 500 most relevant results were analyzed. In total, 5,500 papers were selected for the initial filtering process.
@@
During the initial filtering process, one of the authors manually reviewed the titles of all 5,500 papers. Any paper with a title not related to key terms such as API, security, secret detection, or privacy leakage was excluded. This manual screening step resulted in the selection of 78 papers for further analysis. Subsequently, we conducted a more thorough review of these 78 papers to identify those most relevant to our subject. In this stage, two authors read the abstracts (and, if necessary, the full content) of each paper to determine its relevance. If there was any disagreement, a discussion was held to reach a consensus. For example, the paper “Your Code Secret Belongs to Me: Neural Code Completion Tools Can Memorize Hard-Coded Credentials” secrectllm was initially included after the title-based filtering, but later excluded during this stage as it focuses on credential memorization in neural code completion tools rather than on detection techniques. In total, 19 papers were selected after this step.
@@
We also conducted a snowballing process, a strategy that is commonly adopted in other literature review research Liu_2022. This ensures we do not miss any important papers on the checked-in secret. For each paper that we initially selected, we reviewed references from the selected papers as well as papers that cited them. This process continued until no new relevant papers were identified.
Figure 1: Literature Review Process.
@@
Eventually, we found 29 academic papers that are related to checked-in secret detection. The literature review process is illustrated in Figure 1.
@@
For community tools, we used the six community tools selected by S. K. Basak et al.  emstudy as they conducted a thorough selection process to perform an analysis of checked-in secret detection tools on the open-source repositories.
3.2 RQ1-Result
@@
Among the 29 papers, seven of them developed a checked-in secret detection tool or methodology of our interest, while others are either empirical studies on checked-in secret issue and management or only specific to platforms such as DockerHub. Overall, a total of 13 checked-in secrets detection tools, when combined with community-developed tools, were collected.
@@
We categorized the checked-in secret detection tools into three groups: Intrinsic Value Analysis, Static Analysis, and Machine-Learning-Based Analysis according to their detection approach. Table 2 presents the list of description of all the identified checked-in secret detection tools and their corresponding categorization. Overall, there are seven Intrinsic value analysis tools, one static analysis tool, and five Machine Learning based analysis tools.
Table 2: Categorization, Availability, and Description of Tools
Category Tool Selected Availability Description Intrinsic Value Analysis Meli et al. [7] A Three-Layer Filter combining an entropy-based filter, a word (dictionary) filter, and a pattern filter Gitleak [11] Detects hard-coded secrets in GitHub through regex-based pattern matching Repo Supervisor [12] Detects hard-coded secrets in supported file types via regex-based pattern matching and entropy calculation Trufflehog [9] Detects hard-coded secrets in GitHub via regex-based pattern matching and entropy calculation Whispers [13] Parses structured files (e.g., JSON, YAML, XML) to extract key–value pairs, then applies regex-based rules GGshield [10] Detects hard-coded secrets in GitHub through regex-based pattern matching and entropy calculation GitHub Secret Scanner [14] Detects hard-coded secrets in GitHub through regex-based pattern matching and entropy calculation Static Analysis LeakScope [18] A static analysis tool for Android apps that performs inter-procedural backward data-flow analysis from predefined cloud API calls to extract secrets Machine Learning Analysis PassFinder [8] Identifies secrets in code by analyzing intrinsic string characteristics and surrounding code context Secrethunter [29] Uses reinforcement learning to filter unlikely files on GitHub, then applies regex, entropy checks, and dictionary-based filtering Seagull [30] Detects secrets by analyzing code snippets and their data flow Saha et al. [31] Uses regex-based candidate extraction followed by analyzing string features, entropy, and file metadata Lounici et al. [32] Combines regex-based secret scanning with two ML classifiers—one for file path context and another for code snippet content
@@
Intrinsic value analysis. This type of checked-in secret detection tools focuses on the intrinsic value of the strings under analysis. The analysis is generally done using two processes: regular expression matching and entropy calculation. For regular expression matching, in the cases of certain cloud providers, the generated checked-in secrets often adhere to the same format. For instance, Google’s checked-in secrets always take the form of “ALza******”. Utilizing regular expression matching enables the extraction of such checked-in secrets from the code. For entropy calculation, as checked-in secrets are all randomly generated strings, the randomness of the characters in a string will be a good indicator of checked-in secret. These techniques leverage Shannon entropy to give a good estimation of randomness. A randomly generated string will generally have higher entropy than a normal string.
@@
Static analysis. To the best of our knowledge, LeakScopeLeakScope is the only one tool that extracts checked-in secrets through static analysis . By providing a location where a checked-in secret is used, static analysis can calculate the value of the checked-in secret through data and control flow analysis.
@@
Machine Learning based analysis. This type of checked-in secret detection tool takes into consideration the fact that the contextual surroundings of checked-in secrets follow some specific patterns. Factors such as file type, class name, variable name, and code snippet could be used as features to train a machine learning model to predict whether a string is a checked-in secret.
3.3 Tool Selection for Further Evaluation
The checked-in secrets detection tools within the same category often adopt similar techniques. For example, checked-in secret detection tools based on intrinsic value analysis typically use regular expression matching, entropy calculation, or a combination of both. Since our goal is to analyze the detection approaches rather than individual tools, we chose one representative tool from each category for further evaluation in RQ2 and RQ3. Tool selection was guided not only by effectiveness and adaptability, but also by how well each tool represents its respective approach. We made sure that all the selected tools cover key technologies for each category.
For intrinsic value analysis, we selected the Three-Layer Filter technique proposed by Meli et al. meli because it comprehensively incorporates all major techniques commonly used in this category. Specifically, (1) intrinsic value analysis typically relies on manually crafted regular expressions, and the Three-Layer Filter provides a well-defined and precise set of such patterns; and (2) in addition to regex-based matching, it also applies entropy filtering, word filtering, and pattern filtering, thereby covering the full range of strategies adopted by other tools in this class. Furthermore, (3) it has been shown to be one of the most effective tools in this category emstudy; meli.